What is the limit of (x+sinx+1)/(x+cosx) as x approaches infinity? 

Grade:10

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
^{\lim_{x\rightarrow\infty}}\left ( \frac{x+\sin x+1}{x+\cos x} \right )=^{\lim_{x\rightarrow\infty}}\left ( \frac{1+\cos x}{1+\sin x} \right )=(-)^{\lim_{x\rightarrow\infty}}\left ( \frac{\sin x}{\cos x} \right )= \text{undefined}Regards